### v1

# Multi-scale Autoregressive Models are Laplacian, Discrete, and Latent Diffusion Models in Disguise

Abstract
We revisit Visual Autoregressive (VAR) models through the lens of an iterative-refinement framework. Rather than viewing VAR solely as next-scale autoregression, we formalise it as a deterministic forward process that constructs a Laplacian-style latent pyramid, paired with a learned backward process that reconstructs it in a small number of coarse-to-fine steps. This view connects VAR to denoising diffusion and isolates three design choices that help explain its efficiency and fidelity: refining in a learned latent space, casting prediction as discrete classification over code indices, and partitioning the task by spatial frequency. We run controlled experiments to quantify each factor’s contribution to fidelity and speed, and we outline how the same framework extends to permutation-invariant graph generation and to probabilistic, ensemble-style medium-range weather forecasting. The framework also suggests practical interfaces for VAR to leverage tools from the diffusion ecosystem while retaining few-step, scale-parallel generation.
1 Introduction
Generative modelling spans two complementary paradigms with distinct strengths and limitations [1]. Next-token autoregressive (AR) models fit discrete sequence generation (e.g., text) efficiently [2], but struggle with high-fidelity image generation because their imposed causal orders, such as raster-scan, disrupt natural spatial structures [3, 4]. Diffusion models (DDPM) [5], on the other hand, excel on continuous data, such as images and audio [6]. However, they typically require hundreds to thousands of refinement steps [5] and requires adjustments for discrete outputs [7].
Visual Autoregressive (VAR) models [4] were proposed as a coarse-to-fine alternative that narrows the fidelity gap between diffusion and AR. VAR generates images by predicting the next higher-resolution scale conditioned on all lower-resolution scales, and producing every token at a given scale in one pass rather than token-by-token. In practice, this scale-parallel strategy achieves competitive image quality with significantly fewer iterations, rendering it faster than diffusion.
VAR attributes its gains primarily to the next-scale prediction framework [4]. However, scale-wise refinement alone is not new: the classic Laplacian pyramid [8] follows the same coarse-to-fine idea, and several diffusion models already explore multi-scale structures [9, 10, 11, 12], yet they arguably did not achieve the same fidelity and speed as VAR. It is therefore not immediately clear how scale refinement is the main reason for VAR’s success. This observation motivates a further analysis to explain such attributions.
Figure 1: Iterative refinement in two paradigms. VAR refines by predicting next-scale discrete codes across a small number of scales, while diffusion refines by denoising continuous latents over multiple timesteps.
Recent work advances along two tracks: incremental VAR modifications and broader developments in diffusion. On the VAR side, most studies address engineering bottlenecks: reducing KV-cache and long-sequence overhead with collaborative decoding [13], localising scale refinement to first-order Markov to boost efficiency [14], and reducing codebook bottlenecks by introducing an infinite-vocabulary classifier with bitwise self-correction [15]. On the diffusion side, methods aim to reduce the step count or per-step cost by employing one- or few-step consistency objectives [16], progressive distillation [17], and fast ODE solvers [18], often in conjunction with latent-consistency adapters. Because these efforts proceed in parallel rather than jointly, a unified iterative-refinement perspective would not only allow for consolidation in research efforts but also open paths for VAR to leverage the diffusion ecosystem.
Core contributions.
In this work, we derive VAR as an iterative refinement model rather than a purely multi-scale autoregressive model, enabling a direct connection to denoising diffusion.
We address two questions:
  1. 1.
    In the iterative-refinement formulation of VAR, how can we decompose and quantify which elements drive its success?
  2. 2.
    What implications does this insight have for generative tasks beyond images, including molecular graph generation and weather forecasting?
2 Background
2.1 Visual Autoregressive Models
VQ-VAE tokeniser. VAR uses a vector-quantized autoencoder that turns an input image $x\in\mathbb{R}^{H\times W\times 3}$ into a grid of discrete codes. An encoder $E_{\psi}$ produces a continuous feature map $z_{e}=E_{\psi}(x)$. Each vector in $z_{e}$ is quantized to its nearest codebook embedding $e_{i}$ from a learnable codebook $\mathcal{E}=\{e_{i}\}_{i=1}^{V}$, yielding the quantized map $z_{q}$ that is fed to a decoder $D_{\phi}$ to reconstruct $\hat{x}=D_{\phi}(z_{q})$. Training uses the VQ-VAE objective with a reconstruction term and the codebook/commitment terms:
$\mathcal{L}_{\mathrm{VQ}}\;=\;\underbrace{\|x-\hat{x}\|_{2}^{2}}_{\text{reconstruction}}\;+\;\underbrace{\|\,\mathrm{sg}[z_{e}]-e\,\|_{2}^{2}}_{\text{codebook}}\;+\;\beta\,\underbrace{\|\,z_{e}-\mathrm{sg}[e]\,\|_{2}^{2}}_{\text{commitment}}.$ (1)
where $\mathrm{sg}[\cdot]$ denotes the stop-gradient operator. This loss causes the encoder to commit to discrete codes, while the codebook tracks encoder outputs via vector quantisation, providing discrete token grids suitable for autoregression. VAR replaces a single-scale grid with a stack of token maps $(r^{(1)},\dots,r^{(S)})$ at increasing resolutions produced by a multi-scale VQ tokeniser.
Next-scale autoregression. Let $r^{(k)}$ denote the token map at scale $k$. VAR models the joint as
$p(r^{(1)},\ldots,r^{(S)})\;=\;\prod_{k=1}^{S}p\!\left(r^{(k)}\mid r^{(<k)}\right),$ (2)
and predicts all tokens within a scale in parallel, using attention that is conditioned only on the prefix lower-resolution maps $r^{(\leq k-1)}$ (block-causal across scales). This shift from next-token to next-scale prediction preserves 2D locality and reduces generation complexity relative to raster AR, as each refinement step operates on a 2D grid at a single scale.
It is important to note that the tokeniser and generator operate in residual scales: at level $k$ the encoder forms a residual token map by subtracting the upsampled contribution of coarser scales ($x\!\leftarrow\!x-\phi_{k}(z_{k})$), and at generation time the model predicts residuals and reconstructs by upsample-and-add across all scales ($\hat{x}\!\leftarrow\!\hat{x}+\phi_{k}(z_{k})$). [4]
2.2 Denoising Diffusion Probabilistic Models
DDPMs [5] define a fixed forward (noising) Markov chain
$q(x_{t}\mid x_{t-1})=\mathcal{N}\!\big(\sqrt{1-\beta_{t}}\,x_{t-1},\,\beta_{t}I\big),\quad t=1,\dots,T,$ (3)
which admits a closed-form marginal
$q(x_{t}\mid x_{0})=\mathcal{N}\!\big(\sqrt{\bar{\alpha}_{t}}\,x_{0},\,(1-\bar{\alpha}_{t})I\big),\quad\bar{\alpha}_{t}=\textstyle\prod_{s=1}^{t}\alpha_{s},\;\alpha_{t}=1-\beta_{t}.$ (4)
The reverse chain is learned as Gaussian conditionals
$p_{\theta}(x_{t-1}\mid x_{t})=\mathcal{N}\!\big(\mu_{\theta}(x_{t},t),\,\Sigma_{\theta}(t)\big),$ (5)
by optimizing a variational bound that decomposes into tractable KL terms between Gaussians (thanks to the closed-form forward posterior $q(x_{t-1}\mid x_{t},x_{0})$). In practice, parameterizing $\mu_{\theta}$ via $\varepsilon_{\theta}$ (predicting the added noise) yields the simplified training step
$\min_{\theta}\;\mathbb{E}_{t,x_{0},\varepsilon}\,\big\|\,\varepsilon-\varepsilon_{\theta}\!\big(\sqrt{\bar{\alpha}_{t}}\,x_{0}+\sqrt{1-\bar{\alpha}_{t}}\,\varepsilon,\;t\big)\big\|_{2}^{2},$ (6)
and sampling iterates $x_{T}\sim\mathcal{N}(0,I)\rightarrow x_{0}$ using the learned reverse conditionals.
There are two important remarks about DDPM: (i) Closed-form marginals and posteriors make training low-variance and fully supervised at each noise level (the KL decomposition compares Gaussians in closed form). (ii) Many-step generation: quality improves with the number of reverse steps $T$, but naïve DDPM sampling is slow; learning (or interpolating) reverse variances enables strong quality with far fewer steps (tens instead of hundreds), materially improving practicality.
2.3 Laplacian pyramid
The Laplacian pyramid [8] represents an image as a low-pass residual plus a stack of band-pass detail layers arranged from coarse to fine. Let $G_{0}=x$ be the original image and let $\text{down}(\cdot)$ denote low-pass filtering followed by subsampling. The Gaussian pyramid is built by
$G_{k+1}\;=\;\text{down}(G_{k}),\qquad k=0,\dots,S-1.$ (7)
Define $\text{up}(\cdot)$ as expand (upsample) followed by low-pass filtering that is matched to the analysis filter. The Laplacian (detail) levels are then
$L_{k}\;=\;G_{k}\;-\;\text{up}(G_{k+1}),\qquad k=0,\dots,S-1,$ (8)
together with the coarsest residual $G_{S}$. Perfect reconstruction follows by upsample-and-add:
$\widehat{G}_{S}\;=\;G_{S},\qquad\widehat{G}_{k}\;=\;L_{k}\;+\;\text{up}(\widehat{G}_{k+1})\ \ \text{for }k=S-1,\dots,0,$ (9)
yielding $\widehat{x}=\widehat{G}_{0}=x$ when the synthesis operator $\text{up}(\cdot)$ is the adjoint (or appropriately matched) to the analysis operator in $\text{down}(\cdot)$. Each $L_{k}$ behaves as a spatially localised band-pass component, while $G_{S}$ captures the low frequencies.
3 Multi-scale Autoregressive Models as Iterative Refinement Models
3.1 Iterative refinement
We view iterative refinement as generating by a short sequence of learned updates on a state $x$. From an easy initialiser $x_{0}\sim\pi_{0}$, apply step-specific maps to approach the target:
$x_{t+1}=\Phi_{\theta}^{(t)}(x_{t},c),\qquad t=0,\dots,T-1,\quad\hat{x}=x_{T}.$ (10)
Each step is trained to make a local improvement, e.g. by decreasing a task energy or satisfying a consistency relation:
$E(x_{t+1};c)\leq E(x_{t};c)-\delta_{t}\quad\text{or}\quad\mathcal{C}(x_{t+1},x_{t},c)\approx 0.$ (11)
A coarse-to-fine schedule arises by constraining $\Phi_{\theta}^{(t)}$, or the representation, such that early steps set global structure while later steps add detail; this applies to discrete or continuous $x$; in pixel, latent, or other spaces.
Figure 2: Forward process of VAR. The image is encoded into a latent grid, progressively downsampled to build a pyramid. At each level, we upsample the coarser level to compute a residual, then quantise that residual into discrete indices, and also quantise the coarsest level as base codes. This forward path is deterministic and provides the training targets for the generator.
3.2 Rewriting VAR with iterative refinement
We now define VAR as an iterative refinement framework. We work in the discrete latent space produced by a VQ-VAE tokeniser: the encoder maps $x$ to a continuous feature $z_{e}$, which is vector-quantised to an index grid $z^{(0)}$, and the decoder maps quantised latents back to pixels. Let $L^{(0)}\in\mathbb{R}^{H_{0}\times W_{0}\times d}$ denote the corresponding $d$-channel latent obtained by embedding each index into $\mathbb{R}^{d}$. The multi-scale pyramid is then built by deterministic downscaling of latents, $L^{(k)}=\mathcal{S}_{k}\!\big(L^{(k-1)}\big)$, and, in the spirit of a Laplacian pyramid, we represent the hierarchy in residuals $R^{(k-1)}=L^{(k-1)}-\mathcal{U}_{k}\!\big(L^{(k)}\big)$. At each level, these residuals are quantised with per-scale codebooks whose entries lie in $\mathbb{R}^{d}$, and the coarsest latent $L^{(S)}$ is likewise quantised to provide the base of the refinement process.
The forward process.
As in Figure 2 first encode the image into latent space,
$L^{(0)}\;=\;F(x).$ (12)
We then form a latent Gaussian pyramid by deterministic downscaling,
$L^{(k)}\;=\;\mathcal{S}_{k}\!\bigl(L^{(k-1)}\bigr),\qquad k=1,\dots,S,$ (13)
and define latent residuals in Laplacian form,
$R^{(k-1)}\;=\;L^{(k-1)}\;-\;\mathcal{U}_{k}\!\bigl(L^{(k)}\bigr),\qquad k=1,\dots,S.$ (14)
Each residual is pointwise quantised with a $d$-channel codebook, while the coarsest latent may be quantised as the base:
$r^{(k-1)}\;=\;Q^{\mathrm{res}}_{k}\!\bigl(R^{(k-1)}\bigr),\quad k=1,\dots,S,\qquad b^{(S)}\;=\;Q^{\mathrm{base}}_{S}\!\bigl(L^{(S)}\bigr).$ (15)
The forward distribution is therefore fully deterministic,
$q\!\left(b^{(S)},r^{(0)},\dots,r^{(S-1)}\mid x\right)\;=\;\delta\!\bigl(b^{(S)}-Q^{\mathrm{base}}_{S}(L^{(S)})\bigr)\;\prod_{k=1}^{S}\,\delta\!\bigl(r^{(k-1)}-Q^{\mathrm{res}}_{k}(R^{(k-1)})\bigr),$ (16)
The backward process.
Generation proceeds from coarse to fine in latent space by predicting residual indices in parallel at each scale and reconstructing latents via upsample-and-add. Let the $d$-channel embeddings associated with the residual and base codebooks be $\mathrm{emb}^{\mathrm{res}}_{k}:\{0,\dots,V_{k}-1\}\to\mathbb{R}^{d}$ and $\mathrm{emb}^{\mathrm{base}}_{S}:\{0,\dots,V_{S}-1\}\to\mathbb{R}^{d}$. We first sample a factorised prior over the coarsest base codes and embed them,
$b^{(S)}\sim p_{\theta}\!\bigl(b^{(S)}\bigr),\qquad L^{(S)}\;=\;\mathrm{emb}^{\mathrm{base}}_{S}\!\bigl(b^{(S)}\bigr).$ (17)
For $k=S,\dots,1$ we form the upscaled context and a Transformer $\psi_{\theta}$ predicts the entire residual index map at the next finer scale in one shot,
$C^{(k-1)}\;=\;\mathcal{U}_{k}\!\bigl(L^{(k)}\bigr),\qquad p_{\theta}\!\bigl(r^{(k-1)}\mid L^{(k)}\bigr)\;=\;\prod_{i=1}^{H_{k-1}W_{k-1}}\mathrm{Cat}\!\Bigl(r^{(k-1)}_{i}\,;\,\psi_{\theta}\!\bigl(C^{(k-1)}\bigr)_{i,\cdot}\Bigr),$ (18)
followed by residual embedding and latent reconstruction,
$\widehat{R}^{(k-1)}\;=\;\mathrm{emb}^{\mathrm{res}}_{k}\!\bigl(r^{(k-1)}\bigr),\qquad L^{(k-1)}\;=\;C^{(k-1)}\;+\;\widehat{R}^{(k-1)}.$ (19)
The finest latent is finally decoded to pixels,
$\hat{x}\;=\;G\!\bigl(L^{(0)}\bigr).$ (20)
Remark.
In this formulation, (i) tokens within each scale are generated fully in parallel: conditioned on the upsampled context $C^{(k-1)}$, the categorical factors across spatial sites are independent, so the entire residual map $r^{(k-1)}$ is produced in one shot. (ii) The number of refinement steps is small: just $S$ scales (typically S=8S{=}8 in the original setup), versus the hundreds to a thousand time steps used by diffusion samplers. (iii) The (latent) likelihood is exact and tractable because the model defines a finite autoregressive factorization across scales, whereas DDPMs posit a first-order Markov chain $p_{\theta}(x_{0{:}T})=p(x_{T})\prod_{t=1}^{T}p_{\theta}(x_{t-1}\!\mid x_{t})$ and are trained/evaluated via a variational bound (ELBO) built from Gaussian KL terms rather than exact token-level likelihood.
3.3 Discussion
Given this view, both DDPM and VAR are instances of iterative refinement; this perspective surfaces three design choices that may explain why VAR performs strongly relative to DDPM:
Latent space.
Refining in a learned latent space may reduce spatial dimensionality and attenuate pixel-level redundancy, so updates operate on more aggregated features and reduce task complexity. This could improve the signal-to-noise ratio of supervision and shorten the effective sequence length, broadly in line with observations from latent diffusion [19].
Discrete domain.
Framing prediction over categorical indices can rephrase a continuous regression task as classification, which may yield steadier gradients and mitigate regression-to-the-mean effects [20, 21]. A finite codebook also acts as a mild prior over reusable visual primitives and might enable simple stepwise error-correction, though DDPM’s continuous noise targets [5] offer different benefits.
Frequency refinement.
Replacing denoising with next-scale upscaling explicitly partitions the image into $S$ bands of increasing spatial frequency, akin to a Laplacian pyramid [8]. Each refinement step can specialise: early steps set low-frequency layout and colour; later steps add higher-frequency texture and edges. This frequency factorisation may reduce interference present at intermediate DDPM noise levels, where a single update mixes multiple frequencies. It may also align supervision with perceptual importance per band and enable within-scale parallel token generation, thereby improving the fidelity-speed trade-off, even though $S$ must still be tuned.
4 Experiments
We design three small MLP experiments on MNIST [22] to probe the three hypotheses from Section 3.3: (i) refining in a compact latent may ease optimisation, (ii) discrete prediction may stabilise training relative to continuous regression, and (iii) coarse-to-fine banded refinement may reduce interference between frequencies. We maintain models and budgets in alignment and report the mean $\pm$ standard deviation across three seeds.
Common setup.
We use MNIST $28{\times}28$ images normalised to $[0,1]$. We apply additive Gaussian noise with $\sigma{=}0.20$ at train and test unless stated otherwise. All predictors are 2-layer MLPs (width 512, GELU, with LayerNorm applied before the output layer), trained with AdamW (learning rate $10^{-3}$, weight decay $10^{-4}$) for 30 epochs with a batch size of 256. For latent-space experiments, we pretrain a linear autoencoder with a d=64d{=}64 bottleneck using MSE and then freeze it; we denote its encoder and decoder by $E$ and $D$. For Laplacian experiments, we use a fixed separable $[1,4,6,4,1]/16$ blur for analysis.
Setup A: Pixel vs. latent denoising
[…]
Metrics. We report test MSE $\downarrow$, PSNR $\uparrow$, epochs-to-target (MSE $\leq 0.018$) $\downarrow$, and robustness under a noise shift when testing at $\sigma{=}0.30$.
Table 1: Setup A: Pixel vs. latent denoising on MNIST.
Method MSE @ $\sigma{=}0.20$ $\downarrow$ PSNR (dB) $\uparrow$ Epochs to $0.018$ $\downarrow$ MSE @ $\sigma{=}0.30$ $\downarrow$
Pixel-MLP $0.0195{\pm}0.0003$ $17.6{\pm}0.1$ $14{\pm}1$ $0.0280{\pm}0.0010$
Latent-MLP $\mathbf{0.0158}{\pm}0.0002$ $\mathbf{18.9}{\pm}0.1$ $\mathbf{9}{\pm}1$ $\mathbf{0.0221}{\pm}0.0010$
Observation. The latent model reaches a given error in fewer epochs and maintains lower error under a noise shift. This pattern is consistent with the idea that a compact representation may improve signal-to-noise and shorten the effective optimisation path.
Setup B: Discrete classification vs. continuous regression
[…]
Metrics. We report dequantised test MSE $\downarrow$, mini-batch gradient-norm variance (arbitrary units, lower is steadier) $\downarrow$, and relative training time per epoch.
Table 2: Setup B: Discrete prediction stabilises training at fixed capacity.
Head Dequantised MSE $\downarrow$ Grad var ($\times 10^{-3}$) $\downarrow$ Time/epoch (rel.)
Regression (MSE) $0.0204{\pm}0.0004$ $2.4{\pm}0.2$ $1.00$
Softmax-64 (CE) $\mathbf{0.0181}{\pm}0.0003$ $1.5{\pm}0.1$ $1.02$
Bitwise-8 (BCE) $0.0185{\pm}0.0003$ $\mathbf{1.1}{\pm}0.1$ $\mathbf{0.92}$
Observation. Classification-based targets yield steadier gradients and lower reconstruction error at the same budget. The bitwise factorisation further reduces per-step cost and variance. These results suggest that reframing regression as classification can simplify optimisation.
Setup C: Coarse-to-fine residual learning vs. single-shot
[…]
Metrics. We evaluate test MSE $\downarrow$, epochs-to-target (MSE $\leq 0.018$) $\downarrow$, and high-frequency PSNR (HF-PSNR on $R_{0}$) $\uparrow$.
Table 3: Setup C: Coarse-to-fine specialisation reduces interference.
Method MSE $\downarrow$ Epochs to $0.018$ $\downarrow$ HF-PSNR (dB) $\uparrow$
Single-shot (one MLP) $0.0193{\pm}0.0003$ $12{\pm}1$ $16.8{\pm}0.2$
[…]
Depth sweep. Increasing the number of bands beyond two yields diminishing returns:
Table 4: Effect of refinement depth $S$ with matched total parameters.
$S$ (bands) MSE $\downarrow$ Train time/epoch (rel.)
1 $0.0193$ $1.00$
2 $\mathbf{0.0164}$ $1.15$
3 $0.0162$ $1.22$
Observation. Two banded stages already capture most of the gain, while additional splits offer only small improvements at higher cost. This outcome aligns with the idea that a small number of coarse-to-fine refinements can be sufficient in practice.
5 Beyond Images: Applications of Iterative Refinement
5.1 Graph generation
Graph generation [23, 24] seeks to learn distributions over graphs for applications in chemistry, biology, and network science, where the target law should be invariant to node permutations.
A central challenge is ensuring permutation invariance of the induced distribution. Many autoregressive constructions impose an arbitrary node ordering, which can cause the likelihood or sampling trajectory to depend on that ordering and therefore break invariance [25]. Typical remedies include using exchangeable priors and permutation-equivariant architectures (e.g., GNNs where message passing and pooling commute with node permutations), but care is still required in training objectives and decoding to avoid reintroducing order dependence.
Recent work [26] has instantiated a VAR-style, coarse-to-fine generator for graphs and reported competitive sample quality, along with significant inference speedups, suggesting that next-scale prediction may transfer beyond images.
Diffusion-based graph models, on the other hand, establish that, if the score/denoising network is permutation equivariant, the resulting generative distribution is permutation invariant [27].
Why the iterative refinement perspective may help.
Using the iterative-refinement perspective, we can reach an analogous conclusion for VAR. First, VAR generates by a short sequence of refinement steps, mirroring diffusion’s iterative updates. Second, if the refinement map at each scale is a Transformer that omits order-dependent positional encodings, then each step is order-invariant and, thus, permutation-invariant. With an exchangeable base prior at the coarsest scale, the composition of permutation-invariant steps yields a permutation-equivariant distribution over graphs. While this is an informal argument, it provides a concrete design rule: use exchangeable priors and set-structured Transformers (or equivariant GNNs) for all refinement maps to ensure permutation invariance.
5.2 Medium-range weather forecasting
Medium-range (3–15 day) forecasts inform public safety and energy planning, where distribution-aware evaluation matters. GenCast [28] established a skilful, probabilistic ML forecaster by framing the task as generative modelling with diffusion, providing calibrated ensemble trajectories rather than a single path. In contrast, leading deterministic ML forecasters (e.g., GraphCast [29], Pangu [30]) are single-pass generators trained for point predictions and do not provide distributional forecasts.
Why the iterative refinement perspective may help.
An iterative-refinement (VAR-style) forecaster is probabilistic by construction: each scale predicts categorical distributions over discrete latents, so diverse ensemble members arise with low incremental cost. Coarse-to-fine frequency updates also align with atmospheric structure: early scales capture planetary/synoptic patterns while later scales add mesoscale detail, and replace many diffusion timesteps with a small number of scale transitions, thereby improving efficiency without abandoning distributional modelling.
6 Limitations and Future Work
Scale of experiments.
Our empirical study is intentionally small: we use controlled MLP baselines to isolate effects under limited compute. A more complete assessment would train a reference VAR model end to end on standard image benchmarks and compare to its diffusion analogue under similar settings. Specifically: (i) latent vs. pixel: run next-scale refinement in VQ-VAE latents and, as a “no-latent” control, in a Laplacian pixel pyramid; (ii) discrete vs. continuous: keep the pyramid but replace the categorical head with a Gaussian regression head to test quantisation; (iii) multi-scale vs. single-scale: collapse to a single latent grid to assess the value of coarse to fine. For each setting, attach a diffusion head that predicts $\varepsilon$ with $T$ steps in the same space (pixel diffusion, latent diffusion, and discrete diffusion, respectively). Report quality–cost frontiers (e.g., FID or PSNR versus wall-clock and versus network evaluations), calibration (e.g., bits/dim with standard dequantization), and sensitivity to $S$ scales versus $T$ steps.
Deterministic forward process.
In vanilla VAR, the forward path that builds the latent/residual pyramid is fully deterministic. Stochasticity appears only in bitwise self-correction variants (e.g., [15]) that inject Bernoulli noise at the code level. Whether mild forward randomness helps remains an open question and requires further experiments.
Leverage the diffusion ecosystem.
Several techniques from diffusion may transfer directly to VAR’s iterative-refinement view. Examples include classifier-free guidance adapted to categorical heads, self-conditioning across scales, loss re-weighting akin to simplified diffusion objectives, few-step distillation of a large-$S$ refiner into a small-$S$ student, early-exit or schedule search analogous to noise-schedule tuning, and solver-style improvements such as consistency training to reduce the number of refinement transitions. One might also explore rectified-flow style objectives between consecutive scales, latent-consistency adapters for faster sampling, and temperature or top-$k$ calibration per scale to trade fidelity for diversity. Given the fruitful research in the diffusion literature, it is now the time for VAR to leverage many techniques from this field.
7 Conclusion
We reframed Visual Autoregressive models as iterative refiners operating on a latent Laplacian pyramid. This view makes the generation path explicit: a short sequence of coarse-to-fine residual updates, and clarifies how VAR relates to denoising diffusion while retaining exact cross-scale factorisation and within-scale parallelism. Under this lens, three design choices appear central: working in a compact latent space, predicting discrete indices rather than continuous targets, and allocating supervision across frequency bands. Small toy on MNIST support these hypotheses, indicating that a modest number of banded steps can capture most of the benefit.
Beyond images, the same template suggests practical routes for permutation-equivariant graph generation and for efficient probabilistic weather forecasting, where calibrated ensembles are required. The formulation also exposes straightforward interfaces to the diffusion ecosystem, including guidance, consistency-style objectives, and few-step distillation adapted to categorical heads and scale schedules. A fuller assessment will require end-to-end studies with matched backbones on standard benchmarks, but the iterative-refinement framing already offers a common language for analysing, ablating, and improving both VAR and diffusion families.

### v2

@@
 Abstract
-We revisit Visual Autoregressive (VAR) models through the lens of an iterative-refinement framework.
-Rather than viewing VAR solely as next-scale autoregression, we formalise it as a deterministic forward process that constructs a Laplacian-style latent pyramid, paired with a learned backward process that reconstructs it in a small number of coarse-to-fine steps.
-This view connects VAR to denoising diffusion and isolates three design choices that help explain its efficiency and fidelity: refining in a learned latent space, casting prediction as discrete classification over code indices, and partitioning the task by spatial frequency.
-We run controlled experiments to quantify each factor’s contribution to fidelity and speed, and we outline how the same framework extends to permutation-invariant graph generation and to probabilistic, ensemble-style medium-range weather forecasting.
-The framework also suggests practical interfaces for VAR to leverage tools from the diffusion ecosystem while retaining few-step, scale-parallel generation.
+We reinterpret Visual Autoregressive (VAR) models as iterative refinement models to identify which design choices drive their quality–efficiency trade-off.
+Instead of treating VAR only as next-scale autoregression, we formalise it as a deterministic forward process that builds a Laplacian-style latent pyramid, together with a learned backward process that reconstructs samples in a small number of coarse-to-fine steps.
+This formulation makes the link to denoising diffusion explicit and highlights three modelling choices that may underlie VAR’s efficiency and sample quality: refinement in a learned latent space, discrete prediction over code indices, and decomposition by spatial frequency.
+We support this view with controlled experiments that isolate the contribution of each factor to quality and speed.
+We also discuss how the same framework can be adapted to permutation-invariant graph generation and probabilistic medium-range weather forecasting, and how it provides practical points of contact with diffusion methods while preserving few-step, scale-parallel generation.
 1 Introduction
-Generative modelling spans two complementary paradigms with distinct strengths and limitations [1].
-Next-token autoregressive (AR) models fit discrete sequence generation (e.g., text) efficiently [2], but struggle with high-fidelity image generation because their imposed causal orders, such as raster-scan, disrupt natural spatial structures [3, 4].
-Diffusion models (DDPM) [5], on the other hand, excel on continuous data, such as images and audio [6].
-However, they typically require hundreds to thousands of refinement steps [5] and requires adjustments for discrete outputs [7].
-Visual Autoregressive (VAR) models [4] were proposed as a coarse-to-fine alternative that narrows the fidelity gap between diffusion and AR.
-VAR generates images by predicting the next higher-resolution scale conditioned on all lower-resolution scales, and producing every token at a given scale in one pass rather than token-by-token.
-In practice, this scale-parallel strategy achieves competitive image quality with significantly fewer iterations, rendering it faster than diffusion.
-VAR attributes its gains primarily to the next-scale prediction framework [4].
-However, scale-wise refinement alone is not new: the classic Laplacian pyramid [8] follows the same coarse-to-fine idea, and several diffusion models already explore multi-scale structures [9, 10, 11, 12], yet they arguably did not achieve the same fidelity and speed as VAR.
-It is therefore not immediately clear how scale refinement is the main reason for VAR’s success.
-This observation motivates a further analysis to explain such attributions.
+Generative modelling is often framed through two complementary paradigms, each with clear strengths and limitations [27].
+Next-token autoregressive (AR) models are effective for discrete sequence generation (e.g., text) [25], but they tend to struggle in high-fidelity image synthesis because imposed causal orderings, such as raster scan, disrupt natural spatial structure [23, 22].
+Diffusion models (DDPM), by contrast, perform strongly on continuous data such as images and audio [11, 8].
+Their main practical limitation is computational: high-quality sampling often requires hundreds to thousands of refinement steps, and discrete outputs require additional modelling machinery [11, 2].
+Against this backdrop, Visual Autoregressive (VAR) models [22] were introduced as a coarse-to-fine alternative that narrows the fidelity gap between diffusion and AR.
+VAR predicts each higher-resolution scale conditioned on lower-resolution scales, while generating all tokens within a scale in parallel rather than token by token.
+This scale-parallel strategy yields competitive image quality with substantially fewer refinement iterations, which improves inference speed in practice.
+This naturally raises the question of what exactly drives those gains.
+VAR attributes much of its performance to next-scale prediction [22], yet coarse-to-fine refinement itself is not new.
+Classical Laplacian pyramids already use this principle [5], and several diffusion variants also incorporate multi-scale structure [12, 19, 9, 1].
+The key issue, then, is whether scale refinement alone explains VAR’s combined fidelity and speed.
+We therefore focus on a more explicit attribution analysis.
 Figure 1: Iterative refinement in two paradigms.
-VAR refines by predicting next-scale discrete codes across a small number of scales, while diffusion refines by denoising continuous latents over multiple timesteps.
-Recent work advances along two tracks: incremental VAR modifications and broader developments in diffusion.
-On the VAR side, most studies address engineering bottlenecks: reducing KV-cache and long-sequence overhead with collaborative decoding [13], localising scale refinement to first-order Markov to boost efficiency [14], and reducing codebook bottlenecks by introducing an infinite-vocabulary classifier with bitwise self-correction [15].
-On the diffusion side, methods aim to reduce the step count or per-step cost by employing one- or few-step consistency objectives [16], progressive distillation [17], and fast ODE solvers [18], often in conjunction with latent-consistency adapters.
-Because these efforts proceed in parallel rather than jointly, a unified iterative-refinement perspective would not only allow for consolidation in research efforts but also open paths for VAR to leverage the diffusion ecosystem.
+VAR refines by predicting next-scale discrete codes across a small number of scales; each step adds a quantised residual that captures band-limited detail (with finer scales carrying higher-frequency content).
+DDPM refines by denoising continuous latents over many timesteps; at each step, the model estimates and removes an approximately Gaussian noise component.
+A related motivation comes from current research trajectories.
+Recent work has progressed along two largely separate tracks: incremental improvements to VAR and broader acceleration methods for diffusion.
+On the VAR side, studies mainly target engineering bottlenecks, including collaborative decoding for KV-cache and long-sequence overhead [6], first-order Markov localisation of scale refinement for efficiency [29], and infinite-vocabulary classification with bitwise self-correction to ease codebook bottlenecks [10].
+On the diffusion side, effort has focused on reducing either step count or per-step cost through consistency objectives [21], progressive distillation [20], and fast ODE solvers [14], often with latent-consistency adapters.
+A unified iterative-refinement view can connect these lines of work and clarify where methods may transfer between families.
 Core contributions.
-In this work, we derive VAR as an iterative refinement model rather than a purely multi-scale autoregressive model, enabling a direct connection to denoising diffusion.
-We address two questions:
+In this work, we reframe VAR as an iterative refinement model rather than a purely multi-scale autoregressive model. This perspective yields three concrete contributions:
   1. 1.
-    In the iterative-refinement formulation of VAR, how can we decompose and quantify which elements drive its success?
+    A formal connection between VAR and denoising diffusion via deterministic forward decomposition and learned backward refinement in latent space.
   2. 2.
-    What implications does this insight have for generative tasks beyond images, including molecular graph generation and weather forecasting?
+    Controlled ablations that isolate three factors behind VAR’s efficiency-quality trade-off: latent representation, discrete targets, and coarse-to-fine frequency partitioning.
+  3. 3.
+    Design implications for non-image generation settings, with concrete hypotheses for graph generation and probabilistic weather forecasting.
 2 Background
@@
 VQ-VAE tokeniser.
-VAR uses a vector-quantized autoencoder that turns an input image $x\in\mathbb{R}^{H\times W\times 3}$ into a grid of discrete codes.
+VAR uses a vector-quantised autoencoder that turns an input image $x\in\mathbb{R}^{H\times W\times 3}$ into a grid of discrete codes.
 An encoder $E_{\psi}$ produces a continuous feature map $z_{e}=E_{\psi}(x)$.
-Each vector in $z_{e}$ is quantized to its nearest codebook embedding $e_{i}$ from a learnable codebook $\mathcal{E}=\{e_{i}\}_{i=1}^{V}$, yielding the quantized map $z_{q}$ that is fed to a decoder $D_{\phi}$ to reconstruct $\hat{x}=D_{\phi}(z_{q})$.
+Each vector in $z_{e}$ is quantised to its nearest codebook embedding $e_{i}$ from a learnable codebook $\mathcal{E}=\{e_{i}\}_{i=1}^{V}$, yielding the quantised map $z_{q}$ that is fed to a decoder $D_{\phi}$ to reconstruct $\hat{x}=D_{\phi}(z_{q})$.
 Training uses the VQ-VAE objective with a reconstruction term and the codebook/commitment terms:
@@
 by optimizing a variational bound that decomposes into tractable KL terms between Gaussians (thanks to the closed-form forward posterior $q(x_{t-1}\mid x_{t},x_{0})$).
-In practice, parameterizing $\mu_{\theta}$ via $\varepsilon_{\theta}$ (predicting the added noise) yields the simplified training step
+In practice, parameterising $\mu_{\theta}$ via $\varepsilon_{\theta}$ (predicting the added noise) yields the simplified training step
 $\min_{\theta}\;\mathbb{E}_{t,x_{0},\varepsilon}\,\big\|\,\varepsilon-\varepsilon_{\theta}\!\big(\sqrt{\bar{\alpha}_{t}}\,x_{0}+\sqrt{1-\bar{\alpha}_{t}}\,\varepsilon,\;t\big)\big\|_{2}^{2},$ (6)
@@
 A coarse-to-fine schedule arises by constraining $\Phi_{\theta}^{(t)}$, or the representation, such that early steps set global structure while later steps add detail; this applies to discrete or continuous $x$; in pixel, latent, or other spaces.
-Figure 2: Forward process of VAR.
-The image is encoded into a latent grid, progressively downsampled to build a pyramid.
-At each level, we upsample the coarser level to compute a residual, then quantise that residual into discrete indices, and also quantise the coarsest level as base codes.
-This forward path is deterministic and provides the training targets for the generator.
+Figure 2: Deterministic forward process in VAR.
+An input image is encoded into a latent grid, then repeatedly downsampled to form a multi-scale pyramid.
+At each scale, the upsampled coarser latent is subtracted to obtain a residual, which is quantised into discrete code indices; the coarsest latent is quantised as base codes.
+The resulting set of base and residual codes defines deterministic training targets for the backward (generative) process.
 3.2 Rewriting VAR with iterative refinement
 We now define VAR as an iterative refinement framework.
-We work in the discrete latent space produced by a VQ-VAE tokeniser: the encoder maps $x$ to a continuous feature $z_{e}$, which is vector-quantised to an index grid $z^{(0)}$, and the decoder maps quantised latents back to pixels.
-Let $L^{(0)}\in\mathbb{R}^{H_{0}\times W_{0}\times d}$ denote the corresponding $d$-channel latent obtained by embedding each index into $\mathbb{R}^{d}$.
-The multi-scale pyramid is then built by deterministic downscaling of latents, $L^{(k)}=\mathcal{S}_{k}\!\big(L^{(k-1)}\big)$, and, in the spirit of a Laplacian pyramid, we represent the hierarchy in residuals $R^{(k-1)}=L^{(k-1)}-\mathcal{U}_{k}\!\big(L^{(k)}\big)$.
-At each level, these residuals are quantised with per-scale codebooks whose entries lie in $\mathbb{R}^{d}$, and the coarsest latent $L^{(S)}$ is likewise quantised to provide the base of the refinement process.
+We work in the discrete latent space produced by a VQ-VAE tokeniser.
+The encoder maps $x$ to a continuous feature $z_{e}$, which is vector-quantised to an index grid $z^{(0)}$, and the decoder maps quantised latents back to pixels.
+Let $L^{(0)}\in\mathbb{R}^{H_{0}\times W_{0}\times d}$ denote the $d$-channel latent obtained by embedding each index into $\mathbb{R}^{d}$.
+We construct a multi-scale pyramid by deterministic downscaling, $L^{(k)}=\mathcal{S}_{k}\!\big(L^{(k-1)}\big)$.
+In Laplacian form, the hierarchy is represented by residuals $R^{(k-1)}=L^{(k-1)}-\mathcal{U}_{k}\!\big(L^{(k)}\big)$.
+Each residual is quantised with a per-scale codebook in $\mathbb{R}^{d}$, and the coarsest latent $L^{(S)}$ is quantised as the base state.
 The forward process.
@@
 The backward process.
-Generation proceeds from coarse to fine in latent space by predicting residual indices in parallel at each scale and reconstructing latents via upsample-and-add.
+Generation proceeds from coarse to fine in latent space. At each scale, the model predicts residual indices in parallel and reconstructs latents via upsample-and-add.
 Let the $d$-channel embeddings associated with the residual and base codebooks be $\mathrm{emb}^{\mathrm{res}}_{k}:\{0,\dots,V_{k}-1\}\to\mathbb{R}^{d}$ and $\mathrm{emb}^{\mathrm{base}}_{S}:\{0,\dots,V_{S}-1\}\to\mathbb{R}^{d}$.
@@
 Remark.
-In this formulation, (i) tokens within each scale are generated fully in parallel: conditioned on the upsampled context $C^{(k-1)}$, the categorical factors across spatial sites are independent, so the entire residual map $r^{(k-1)}$ is produced in one shot.
-(ii) The number of refinement steps is small: just $S$ scales (typically S=8S{=}8 in the original setup), versus the hundreds to a thousand time steps used by diffusion samplers.
-(iii) The (latent) likelihood is exact and tractable because the model defines a finite autoregressive factorization across scales, whereas DDPMs posit a first-order Markov chain $p_{\theta}(x_{0{:}T})=p(x_{T})\prod_{t=1}^{T}p_{\theta}(x_{t-1}\!\mid x_{t})$ and are trained/evaluated via a variational bound (ELBO) built from Gaussian KL terms rather than exact token-level likelihood.
+In this formulation, (i) tokens within each scale are generated fully in parallel: conditioned on the upsampled context $C^{(k-1)}$, categorical factors across spatial sites are independent, so the entire residual map $r^{(k-1)}$ is produced in one shot.
+(ii) The number of refinement steps is small: just $S$ scales (typically S=8S{=}8 in the original setup), versus hundreds to a thousand time steps in common diffusion samplers.
+(iii) The latent likelihood is exact and tractable because the model defines a finite autoregressive factorisation across scales.
+By contrast, DDPMs posit a first-order Markov chain $p_{\theta}(x_{0{:}T})=p(x_{T})\prod_{t=1}^{T}p_{\theta}(x_{t-1}\!\mid x_{t})$ and are trained/evaluated via an ELBO built from Gaussian KL terms.
 3.3 Discussion
-Given this view, both DDPM and VAR are instances of iterative refinement; this perspective surfaces three design choices that may explain why VAR performs strongly relative to DDPM:
+Given this formulation, DDPM and VAR can both be viewed as iterative refinement methods. Under this shared view, three modeling choices are likely to matter for VAR’s performance relative to DDPM.
 Latent space.
-Refining in a learned latent space may reduce spatial dimensionality and attenuate pixel-level redundancy, so updates operate on more aggregated features and reduce task complexity.
-This could improve the signal-to-noise ratio of supervision and shorten the effective sequence length, broadly in line with observations from latent diffusion [19].
+VAR refines in a learned latent representation rather than raw pixels.
+This typically lowers dimensionality and removes some local redundancy, so each update is applied to more compact, semantically organised features.
+A simpler target space can improve optimisation efficiency and may increase the effective signal-to-noise ratio during training, consistent with observations in latent diffusion [18].
 Discrete domain.
-Framing prediction over categorical indices can rephrase a continuous regression task as classification, which may yield steadier gradients and mitigate regression-to-the-mean effects [20, 21].
-A finite codebook also acts as a mild prior over reusable visual primitives and might enable simple stepwise error-correction, though DDPM’s continuous noise targets [5] offer different benefits.
+VAR predicts categorical code indices, whereas standard DDPM predicts continuous noise.
+Casting prediction as classification can yield more stable gradients than direct regression in some settings [24, 17].
+The finite codebook also imposes a structured representation that can support stepwise correction across scales.
+This does not imply that discrete targets are universally better; continuous noise prediction also has advantages, especially for smooth interpolation and flexible likelihood objectives [11].
 Frequency refinement.
-Replacing denoising with next-scale upscaling explicitly partitions the image into $S$ bands of increasing spatial frequency, akin to a Laplacian pyramid [8].
-Each refinement step can specialise: early steps set low-frequency layout and colour; later steps add higher-frequency texture and edges.
-This frequency factorisation may reduce interference present at intermediate DDPM noise levels, where a single update mixes multiple frequencies.
-It may also align supervision with perceptual importance per band and enable within-scale parallel token generation, thereby improving the fidelity-speed trade-off, even though $S$ must still be tuned.
+VAR’s next-scale updates implement an explicit coarse-to-fine decomposition, similar in spirit to a Laplacian pyramid [5].
+Early stages primarily model low-frequency structure, while later stages focus on high-frequency detail.
+This separation can reduce cross-frequency interference compared with updates that mix multiple frequency regimes in a single step.
+It also enables parallel token prediction within each scale.
+In practice, gains depend on the number of scales $S$, which remains a tunable design choice.
 4 Experiments
-We design three small MLP experiments on MNIST [22] to probe the three hypotheses from Section 3.3: (i) refining in a compact latent may ease optimisation, (ii) discrete prediction may stabilise training relative to continuous regression, and (iii) coarse-to-fine banded refinement may reduce interference between frequencies.
-We maintain models and budgets in alignment and report the mean $\pm$ standard deviation across three seeds.
-Common setup.
-We use MNIST $28{\times}28$ images normalised to $[0,1]$.
-We apply additive Gaussian noise with $\sigma{=}0.20$ at train and test unless stated otherwise.
-All predictors are 2-layer MLPs (width 512, GELU, with LayerNorm applied before the output layer), trained with AdamW (learning rate $10^{-3}$, weight decay $10^{-4}$) for 30 epochs with a batch size of 256.
-For latent-space experiments, we pretrain a linear autoencoder with a d=64d{=}64 bottleneck using MSE and then freeze it; we denote its encoder and decoder by $E$ and $D$.
-For Laplacian experiments, we use a fixed separable $[1,4,6,4,1]/16$ blur for analysis.
-Setup A: Pixel vs. latent denoising
+We run three controlled MNIST [7] studies to probe the hypotheses from Section 3.3: latent-space refinement, discrete targets, and coarse-to-fine frequency decomposition.
+Results are reported as mean $\pm$ standard deviation over three seeds.
+Table 1: Common training protocol across experiments.
+Dataset MNIST ($28\times 28$, normalised to $[0,1]$)
+Train/test corruption Additive Gaussian noise, $\sigma=0.20$ unless stated
+Backbone 2-layer MLP, width 512, GELU, pre-output LayerNorm
+Optimiser AdamW, learning rate $10^{-3}$, weight decay $10^{-4}$
+Training budget 30 epochs, batch size 256
+Latent encoder (when used) Linear autoencoder, bottleneck $d=64$, then frozen
+Pyramid filter (when used) Separable blur $[1,4,6,4,1]/16$
+4.1 Latent vs Pixel Space
 […]
-Metrics. We report test MSE $\downarrow$, PSNR $\uparrow$, epochs-to-target (MSE $\leq 0.018$) $\downarrow$, and robustness under a noise shift when testing at $\sigma{=}0.30$.
+We evaluate test MSE, PSNR, epochs-to-target (MSE $\leq 0.018$), and robustness under a test-time noise shift to $\sigma=0.30$.
 Table 2: Setup A: Pixel vs. latent denoising on MNIST.
@@
 Latent-MLP $\mathbf{0.0158}{\pm}0.0002$ $\mathbf{18.9}{\pm}0.1$ $\mathbf{9}{\pm}1$ $\mathbf{0.0221}{\pm}0.0010$
-Observation.
-The latent model reaches a given error in fewer epochs and maintains lower error under a noise shift.
-This pattern is consistent with the idea that a compact representation may improve signal-to-noise and shorten the effective optimisation path.
-Setup B: Discrete classification vs. continuous regression
-[…]
-Metrics. We report dequantised test MSE $\downarrow$, mini-batch gradient-norm variance (arbitrary units, lower is steadier) $\downarrow$, and relative training time per epoch.
+Latent-MLP reduces MSE from $0.0195$ to $0.0158$, improves PSNR from $17.6$ to $18.9$ dB, reaches the target in 9 epochs instead of 14, and remains better under the $\sigma=0.30$ shift.
+This is consistent with the hypothesis that compact learned representations improve optimization efficiency.
+4.2 Discrete vs Continuous Targets
+We build a $k$-means codebook ($k=64$) on clean latent features $E(x)$, then train a shared MLP trunk with alternative prediction heads.
+The regression head predicts continuous latent vectors with MSE.
+Softmax-64 predicts nearest-code indices with cross-entropy and dequantises to centroids.
+Bitwise-8 predicts a learned 8-bit code with independent Bernoulli heads.
+We report dequantised test MSE, mini-batch gradient-norm variance (lower is steadier), and relative training time per epoch.
 Table 3: Setup B: Discrete prediction stabilises training at fixed capacity.
@@
 Bitwise-8 (BCE) $0.0185{\pm}0.0003$ $\mathbf{1.1}{\pm}0.1$ $\mathbf{0.92}$
-Observation.
-Classification-based targets yield steadier gradients and lower reconstruction error at the same budget.
-The bitwise factorisation further reduces per-step cost and variance.
-These results suggest that reframing regression as classification can simplify optimisation.
-Setup C: Coarse-to-fine residual learning vs. single-shot
-[…]
-Metrics. We evaluate test MSE $\downarrow$, epochs-to-target (MSE $\leq 0.018$) $\downarrow$, and high-frequency PSNR (HF-PSNR on $R_{0}$) $\uparrow$.
+Discrete targets improve both stability and accuracy at fixed capacity.
+Softmax-64 lowers MSE from $0.0204$ to $0.0181$ with lower gradient variance.
+Bitwise-8 further reduces gradient variance and is the fastest per epoch, indicating that factorised discrete targets can improve the quality-cost trade-off.
+4.3 Single-Stage vs Coarse-to-Fine Refinement
+We construct a two-level Laplacian decomposition with low-pass component $G_{1}$ ($14\times 14$) and residual $R_{0}$ ($28\times 28$).
+The single-stage baseline predicts the clean image directly from noisy input.
+The coarse-to-fine variant first denoises $G_{1}$ with $\mathrm{MLP}_{\text{coarse}}$, then predicts $R_{0}$ with $\mathrm{MLP}_{\text{resid}}$ conditioned on the upsampled $\widehat{G}_{1}$ and the noisy input, and reconstructs $\hat{x}=\mathrm{up}(\widehat{G}_{1})+\widehat{R}_{0}$.
+Total parameters are matched to the baseline.
+We evaluate test MSE, epochs-to-target (MSE $\leq 0.018$), and high-frequency PSNR (HF-PSNR) computed on $R_{0}$.
 Table 4: Setup C: Coarse-to-fine specialisation reduces interference.
@@
 […]
-Depth sweep. Increasing the number of bands beyond two yields diminishing returns:
+Increasing the number of bands beyond two yields diminishing returns:
 Table 5: Effect of refinement depth $S$ with matched total parameters.
@@
 3 $0.0162$ $1.22$
-Observation.
-Two banded stages already capture most of the gain, while additional splits offer only small improvements at higher cost.
-This outcome aligns with the idea that a small number of coarse-to-fine refinements can be sufficient in practice.
+The coarse-to-fine model reduces MSE from $0.0193$ to $0.0164$, reaches the target 5 epochs earlier, and improves HF-PSNR by $1.4$ dB.
+The depth sweep shows that most gains appear at two bands, with only marginal improvement at three bands and higher per-epoch cost.
 5 Beyond Images: Applications of Iterative Refinement
@@
 Why the iterative refinement perspective may help.
-Using the iterative-refinement perspective, we can reach an analogous conclusion for VAR.
-First, VAR generates by a short sequence of refinement steps, mirroring diffusion’s iterative updates.
-Second, if the refinement map at each scale is a Transformer that omits order-dependent positional encodings, then each step is order-invariant and, thus, permutation-invariant.
-With an exchangeable base prior at the coarsest scale, the composition of permutation-invariant steps yields a permutation-equivariant distribution over graphs.
-While this is an informal argument, it provides a concrete design rule: use exchangeable priors and set-structured Transformers (or equivariant GNNs) for all refinement maps to ensure permutation invariance.
+The iterative-refinement view suggests a direct design hypothesis for VAR on graphs.
+VAR already generates through a short sequence of refinement steps, analogous to diffusion updates.
+If each refinement map is implemented by a permutation-equivariant architecture (for example, a set-structured Transformer without order-dependent positional encodings or an equivariant GNN), and the coarsest prior is exchangeable, then the resulting generation process is expected to preserve permutation symmetry.
+This is a design principle rather than a formal proof for the full model class.
+Empirical checks of invariance and likelihood consistency across node relabelings are still required.
 5.2 Medium-range weather forecasting
@@
 Why the iterative refinement perspective may help.
-An iterative-refinement (VAR-style) forecaster is probabilistic by construction: each scale predicts categorical distributions over discrete latents, so diverse ensemble members arise with low incremental cost.
-Coarse-to-fine frequency updates also align with atmospheric structure: early scales capture planetary/synoptic patterns while later scales add mesoscale detail, and replace many diffusion timesteps with a small number of scale transitions, thereby improving efficiency without abandoning distributional modelling.
+An iterative-refinement VAR-style forecaster could be probabilistic by construction if each scale predicts distributions over discrete latents.
+In that case, multiple ensemble members can be sampled with low incremental cost per member.
+The coarse-to-fine decomposition also aligns with atmospheric scale separation: early levels represent planetary and synoptic structure, while later levels refine mesoscale detail.
 6 Limitations and Future Work
 Scale of experiments.
-Our empirical study is intentionally small: we use controlled MLP baselines to isolate effects under limited compute.
-A more complete assessment would train a reference VAR model end to end on standard image benchmarks and compare to its diffusion analogue under similar settings.
-Specifically: (i) latent vs.
-pixel: run next-scale refinement in VQ-VAE latents and, as a “no-latent” control, in a Laplacian pixel pyramid; (ii) discrete vs.
-continuous: keep the pyramid but replace the categorical head with a Gaussian regression head to test quantisation; (iii) multi-scale vs.
-single-scale: collapse to a single latent grid to assess the value of coarse to fine.
-For each setting, attach a diffusion head that predicts $\varepsilon$ with $T$ steps in the same space (pixel diffusion, latent diffusion, and discrete diffusion, respectively).
-Report quality–cost frontiers (e.g., FID or PSNR versus wall-clock and versus network evaluations), calibration (e.g., bits/dim with standard dequantization), and sensitivity to $S$ scales versus $T$ steps.
+Our empirical study is intentionally limited in scope.
+We use controlled MLP baselines to isolate specific factors under modest compute budgets.
+A broader assessment should train a full VAR model end-to-end on standard image benchmarks and compare it with a diffusion baseline under matched architectures and training budgets.
+Useful ablations include: (i) latent vs.
+pixel, by running next-scale refinement in VQ-VAE latent space and in a Laplacian pixel pyramid as a no-latent control; (ii) discrete vs.
+continuous, by replacing categorical prediction heads with Gaussian regression heads while keeping the same pyramid structure; and (iii) multi-scale vs.
+single-scale, by collapsing to a single latent grid.
+For each setting, one can add a diffusion objective in the same representation space (pixel, latent, or discrete) and compare quality–cost trade-offs (e.g., FID/PSNR versus wall-clock and function evaluations), calibration (e.g., bits/dim with standard dequantisation), and sensitivity to refinement depth $S$ versus diffusion steps $T$.
 Deterministic forward process.
-In vanilla VAR, the forward path that builds the latent/residual pyramid is fully deterministic.
-Stochasticity appears only in bitwise self-correction variants (e.g., [15]) that inject Bernoulli noise at the code level.
-Whether mild forward randomness helps remains an open question and requires further experiments.
-Leverage the diffusion ecosystem.
-Several techniques from diffusion may transfer directly to VAR’s iterative-refinement view.
-Examples include classifier-free guidance adapted to categorical heads, self-conditioning across scales, loss re-weighting akin to simplified diffusion objectives, few-step distillation of a large-$S$ refiner into a small-$S$ student, early-exit or schedule search analogous to noise-schedule tuning, and solver-style improvements such as consistency training to reduce the number of refinement transitions.
-One might also explore rectified-flow style objectives between consecutive scales, latent-consistency adapters for faster sampling, and temperature or top-$k$ calibration per scale to trade fidelity for diversity.
-Given the fruitful research in the diffusion literature, it is now the time for VAR to leverage many techniques from this field.
+In standard VAR, the forward construction of the latent/residual pyramid is deterministic.
+Stochasticity is introduced only in some variants, such as bitwise self-correction methods that inject Bernoulli noise at the code level [10].
+The effect of adding controlled stochasticity to the forward process is not yet clear and warrants systematic evaluation.
+Leveraging diffusion methods.
+The iterative-refinement view suggests several diffusion-inspired techniques that may be applicable to VAR.
+Candidate directions include classifier-free guidance for categorical heads, self-conditioning across scales, loss reweighting, few-step distillation from larger-$S$ models, and schedule design or early-exit strategies.
+Other possibilities include consistency-style objectives between scales, latent-consistency adapters for faster sampling, and per-scale temperature or top-$k$ calibration to tune fidelity-diversity trade-offs.
+These directions should be treated as hypotheses for future empirical study.
 7 Conclusion
 We reframed Visual Autoregressive models as iterative refiners operating on a latent Laplacian pyramid.
-This view makes the generation path explicit: a short sequence of coarse-to-fine residual updates, and clarifies how VAR relates to denoising diffusion while retaining exact cross-scale factorisation and within-scale parallelism.
+This view makes the generation path explicit as a short sequence of coarse-to-fine residual updates and clarifies how VAR relates to denoising diffusion while retaining exact cross-scale factorisation and within-scale parallelism.
 Under this lens, three design choices appear central: working in a compact latent space, predicting discrete indices rather than continuous targets, and allocating supervision across frequency bands.
-Small toy on MNIST support these hypotheses, indicating that a modest number of banded steps can capture most of the benefit.
+Small MNIST toy experiments support these hypotheses, indicating that a modest number of banded steps can capture most of the benefit.
 Beyond images, the same template suggests practical routes for permutation-equivariant graph generation and for efficient probabilistic weather forecasting, where calibrated ensembles are required.